OpenStudy (anonymous):

Which values of x will make the absolute value equation true? |x – 2| = 6 A. {–4, –8} B. {–4, 8} C. {4, 8} D. no solution

OpenStudy (jdoe0001):

solve for "x" keep in mind that |x – 2| really means +1(x-2) AND -1(x-2) therefore \(\bf |x - 2| = 6 \implies +1(x - 2)=6 1, -1(x - 2)=6 \) so you'd get 2 values for "x"

OpenStudy (anonymous):

-4 , 8

OpenStudy (anonymous):

yea thats correct :)